OpenStudy (anonymous):

Solve the following equation, answer as a reduced, mixed number 3(x - 2) + 4(2x - 6) = 6(x - 4) + 8(2x + 1)

mathslover (mathslover):

\[\huge{3x-6+8x-24=6x-24+16x+8}\] can u solve this now ?

OpenStudy (anonymous):

3x+8x-6-24=6x+16x-24+8 11x-30=22x-16 -14=11x x=-14/11

OpenStudy (anonymous):

see that i understand but when it comes to putting it as a mixed number

OpenStudy (anonymous):

i get lost

OpenStudy (anonymous):

3x-6+8x-24=6x-24+16x+8 11x-6=22x+8 -14=11x x=-14/11

OpenStudy (anonymous):

x=1 -3/11

OpenStudy (anonymous):

put one as whole with -3/11 as a fraction,together they form mixed no.